# LUPET: Incorporating Hierarchical Information Path into Multilingual ASR

### v1

@@ Abstract
Many factors have separately shown their effectiveness on improving multilingual ASR. They include language identity (LID) and phoneme information, language-specific processing modules and cross-lingual self-supervised speech representation, etc. However, few studies work on synergistically combining them to contribute a unified solution, which still remains an open question. To this end, a novel view to incorporate hierarchical information path LUPET into multilingual ASR is proposed. The LUPET is a path encoding multiple information in different granularity from shallow to deep encoder layers. Early information in this path is beneficial for deriving later occurred information. Specifically, the input goes from LID prediction to acoustic unit discovery followed by phoneme sharing, and then dynamically routed by mixture-of-expert for final token recognition. Experiments on 10 languages of Common Voice examined the superior performance of LUPET. Importantly, LUPET significantly boosts the recognition on high-resource languages, thus mitigating the compromised phenomenon towards low-resource languages in a multilingual setting.
Index Terms: Multilingual ASR, language identity, self-supervised speech representation learning, mixture-of-expert
@@ 1 Introduction
Multilingual ASR aims to use one unified model to transcribe speech in multiple languages to their textual representations [1, 2, 3, 4]. This offers two great advantages [5] that (1) the per-language training cost and deployment effort are largely reduced; (2) recognition performance on low-resource languages can be improved a lot, benefiting from the shared language-invariant knowledge. Previous studies [6] showed that directly training a fully-shared end-to-end (E2E) model on a multilingual corpus is a simple yet effective solution (vanilla). The multilingual corpus is formed by mixing corpora of different languages, meanwhile creating a shared vocabulary set. However, due to the heterogeneous nature among languages [7, 6], the vanilla scheme often causes high-resource languages to compromise their performance towards low-resources in multilingual training, i.e., worse than their monolingual counterparts.
To mitigate the degradation on high-resource languages, recent research considers incorporating different multilingual-related information, e.g., language identity (LID) [8, 9, 10, 11] and phoneme [12, 13, 14] or language-specific architecture, e.g., mixture-of-expert (MoE) [15, 16, 17, 18, 19]. Either including the above information or adopting MoE architecture, one common feature is that they are both in a supervised training framework. Self-supervised learning (SSL) has presented numerous successful practices, like wav2vec2.0 [20], HuBERT [21] and XLSR [22], etc, on speech recognition and should not be neglected. Usually, a two-stage training scheme, i.e., pre-train and then fine-tune, is widely applied. In [23], it was found that jointly unsupervised and supervised training (JUST) on multilingual ASR can outperform the two-stage scheme and prevent forgetting knowledge learned by SSL. The SSL in JUST consists of a contrastive loss and a masked language model (MLM) loss to learn discrete units for better contextualized representations.
In this paper, we consider how to synergistically combine the aforementioned useful factors to boost overall performance. A hierarchical information path of multilingual ASR is hypothesized and empirically examined. The proposed path is from LID prediction to acoustic Unit discovery followed by international Phonetic alphabet (IPA) sharing, then routed by MoE for final multilingual Token recognition (debudded as LUPET  [^1] LUPET in Filipino means “Awesome”.). LUPET seamlessly includes multiple pieces of information in different granularity, e.g., LID, acoustic unit, phoneme, and token. Within this path, the information of the shallow layer is beneficial for deriving those occurred in deeper layers, thus being hierarchical. Importantly, these information labels are either straightforward to obtain or derived in an SSL manner.
LUPET can be easily integrated into the vanilla ASR architecture, significantly boosting the multilingual recognition performance. We conduct extensive experiments to evaluate the proposed LUPET. Experiments on 10 languages of the Common Voice [24] show that, compared to the vanilla method, our proposed LUPET gives 19.73% and 12.27% relative averaged word error rate reductions with CTC and attention decoding, respectively. Furthermore, LUPET beats all previous baselines and exhibits superior performance on high-resource languages.
@@
2 LUPET Method
This section starts with a brief introduction of a vanilla E2E multilingual ASR architecture and then illustrates how to incorporate the proposed LUPET information path.
@@ 2.1 Vanilla E2E Multilingual ASR
The E2E multilingual ASR architecture we adopt is the hybrid CTC-Attention conformer [25, 26]. It consists of three components, namely encoder, decoder, and CTC [27] layer. The encoder takes an acoustic feature sequence $\mathbf{X}=\{\mathbf{x}_{t}\}_{t=1}^{T}$ as input and converts it to hidden representation $\mathbf{H}=\{\mathbf{h}_{t}\}_{t=1}^{T^{{}^{\prime}}}$, where $T$ and $T^{{}^{\prime}}$ denote the number of original frames and the sub-sampled frames. The $\mathbf{H}$ is then forwarded to two classification branches for predicting the token sequence $\mathbf{Y}=\{y_{u}\in\mathcal{V}\}_{u=1}^{U}$, where $\mathcal{V}$ is a shared multilingual vocabulary set built by BPE [28] and $U$ denotes the number of tokens. One classification branch, i.e., decoder, conditions $\mathbf{H}$ to autoregressively compute token-level posterior $p(y_{u}|\mathbf{H},y_{1:u-1})$ via a cross attention mechanism. The attention loss is given as
$\mathcal{L}_{Attn}=-\sum_{u=1}^{U}logp(y_{u}|\mathbf{H},y_{1:u-1}).\vspace{-2mm}$ (1)
@@
where $\mathbf{z}_{t}$ is the logits over $\mathcal{V}\cup\emptyset$ and $B^{-1}$ is the inverse function that gives all the valid alignment paths between input sequence $\mathbf{H}$ and output sequence $\mathbf{Y}$. The blank token $\emptyset$ is specially designed by CTC for aligning $\mathbf{H}$ and $\mathbf{Y}$. $\mathbf{Z}=\{\mathbf{z}_{t}\}_{t=1}^{T^{{}^{\prime}}}$ represents one possible alignment path. The training objective of hybrid CTC-Attention is a linear combination of Eq. 1 and Eq. 2:
@@ 2.2 Incorporating LUPET
Recent studies have pointed out (1) Both LID and phoneme information are beneficial for multilingual training [11, 14]; (2) Design language-specific modules to process language-specific information is useful to reduce language interference [17, 19]; (3) The success of self-supervised cross-lingual representation learning applied in ASR [22, 29]. Although these factors' effectiveness has been verified separately, how to synergistically combine them to contribute a better solution from a unified perspective remains as an open question. The proposed LUPET provides a novel view from the multilingual hierarchical information path. From LID to acoustic unit followed by phoneme then go through MoE routing to the final token, the information that occurred in the early position of the path is assumed to contribute to the prediction of later information of the path.
@@
where $LIN$ denotes a linear layer to keep the hidden dimension and $\mathbf{H}^{s^{\prime}}$ is the input representations for the next encoder layer.
The lower-middle layers of encoder $Enc^{lm}$ are utilized to perform acoustic unit discovery. Similar to BEST-RQ [29], a random-projection quantizer including a projection matrix $Proj^{u}$ and codebook $\mathbb{C}$ is applied and none of the parameters are trainable. Vector quantization (VQ) is carried out on acoustic features $\mathbf{X}$ to produce discrete labels
@@
where $Sub$ represents the subsample operation. The output of $\mathbb{C}$ are the indices of the nearest code vectors in the codebook.
@@
To incorporate the hierarchical information, from LID, acoustic unit, phoneme, and token, the ultimate objective function of LUPET is given as a linear combination of Eq. (3, 4, 7, 8):
@@ 3.1 Dataset
The 10 languages, namely English (en), French (fr), Spanish (es), Chinese (zh), Italian (it), Russian (ru), Portuguese (pt), Turkish (tr), Dutch (nl) and Tatar (tt) from the public available Common Voice 13.0 [24] are selected for our multilingual ASR experiments. The language coverage includes high-resource languages, e.g., English with around 2,280 hours of training data, and low-resource languages, e.g., Tatar with only about 20 hours. The detailed training and testing statistics are listed in Tab. 1. Note that zh includes Mandarin, Taiwanese, and Cantonese. A standard text normalization (the same as in Whisper [2]) is applied to all the transcriptions of the dataset.
@@ 3.2.1 Vanilla
The vanilla model adopts a hybrid CTC-Attention architecture. The encoder has 12 conformer layers with 8 attention heads and 512 hidden dimensions, while the decoder has 6 transformer layers [30]. The CTC weight $\lambda$ in Eq. 3 is set to $0.3$. The input acoustic feature to the network is the typical 80-dimensional log-Mel filterbank. The output vocabulary used is derived from Whisper's tokenizer. This tokenizer was obtained by BPE using UTF-8 bytes of the entire training dataset of Whisper.
@@ 3.2.2 LUPET
Compared to vanilla, the encoder architecture has several modifications by incorporating LUPET. The output positions of { $Enc^{s}$, $Enc^{lm}$, $Enc^{um}$, $Enc^{d}$} are at the {3-th,6-th,9-th,12-th} layer of the original encoder, respectively. When MLM loss takes effect, acoustic feature $\mathbf{X}$ is randomly masked consecutive 20 frames with probability $p=0.01$. Codebook $\mathbb{C}$ of the random-projection quantizer has a size of 8192 and a dimension of 16. IPA sequence per-utterance is obtained using an open-sourced toolkit phonemizer  [^2] https://github.com/bootphon/phonemizer. In each layer of $Enc^{d}$, MoE including 8 FFN experts is used to replace the end-FFN of the original conformer layer. In Eq. 10, the weight coefficients $w_{1}$, $w_{2}$, and $w_{3}$ are set to 0.3, 0.07, and 0.3, respectively.
@@ 3.2.3 Baselines
Several baselines are used for comparison: (1) Mono, monolingual ASR with vanilla architecture is trained per-language. (2) Oracle_LID, append the pre-known LID embedding to input acoustic feature for multilingual ASR training. (3) MoE [17], keep the $Enc^{d}$ of LUPET and remove other auxiliary losses, hidden representation $\mathbf{H}^{um}$ is used as the input to the routing network. (4) LID_SC [11], LID prediction by CTC and LID information self-conditioning (SC) are performed. (5) Whisper, whisper-large-v2 with oracle LID is used for decoding.
@@ 3.3 Training Scheme and Evaluation Metric
We implement the vanilla and our proposed LUPET methods on the Wenet toolkit [31]. The model is trained with Adam [32] optimizer with a learning rate (LR) of $31e-$ 3. LR schedule has a warmup step of $15000$. Batch size is set to $12$ with $accum\_grad=16$. 8 V100 GPUs are used for DDP training. Each multilingual model is trained for 50 epochs and each monolingual model is trained for 100 epochs. If not specified otherwise, MLM takes effect from epoch 5 to 30. The final model for decoding is obtained by averaging the $10$ best models with the lowest validation losses. Character error rate (CER) for Chinese and word error rate (WER) for other languages are used to measure the system's performance.
@@ 4 Results and Analysis
Table 2: WER (%) results of different systems on 10 languages of Common Voice by CTC greedy decoding. In the LUPET block, the backslash /  represents the ablation study by removing the following component, where U = acoustic unit discovery, P = IPA sharing, and L = LID prediction. $w_{2}=1$ means the weight coefficient of $\mathcal{L}_{mlm}$ is set as 1. Uto50ep means U takes effect until 50 epochs.
@@ 4.1 Performance Comparison to Monolingual System
A desirable multilingual ASR is expected to present better performance than its monolingual counterparts (Mono). As shown in Fig. 2, LUPET and other baselines are used to compare performance to monolingual systems on the test sets of 10 languages. The x-axis follows an order from high-resource to low-resource languages. The y-axis denotes the relative WER to Mono, where the more negative value represents the lower WER. It is clear to see all curves are basically decreasing except the peak at ru. The decreasing trend is straightforward as the more low-resourced languages can achieve more significant performance gains. Multilingual training brings degradation to the Russian (ru) language in most cases. We speculate this may be due to the compromised phenomenon from Russian (ru) to Tarta (tt). The language tt achieves above 60% relative WER reduction via multilingual training at the cost of side effects on language ru, which shares the same language family as tt.
@@
4.2 LUPET's Effectiveness Verification
Tab. 2 presents the WER results of different systems on 10 languages. Averaged WER are calculated for overall comparison. The top-5 languages, with more training data, are roughly referred to as high-resources, while the remaining five are low-resources. Having similar observation from Fig. 2, LUPET gives significantly better performance on all languages compared to other baselines. The benefits of LID prediction and MoE routing structure have been well verified by system LID_SC and MoE.
@@ 4.2 LUPET's Effectiveness Verification
Disabling both U and P (IPA sharing prediction) exhibits worse results. Tab. 2 provides two comparison views to understand the independent effect of the component P. (1) LUPET / LU can be seen as ``MoE + P". Comparing it with MoE, IPA sharing is found to be beneficial for all languages except tt. (2) When only removing P from LUPET, WERs on all languages basically degrade. Low-resource languages clearly give the worse results, where tr increase the absolute WER above 4%. One possible reason is that the independent U would lead to information loss due to the masking mechanism in MLM, especially for low-resource languages. In LUPET, with the help of intermediate phoneme prediction (P), the loss of information is largely mitigated, thus not presenting much worse results on low-resource languages.

### v3

@@ Abstract
Toward high-performance multilingual automatic speech recognition (ASR), various types of linguistic information and model design have demonstrated their effectiveness independently. They include language identity (LID), phoneme information, language-specific processing modules, and cross-lingual self-supervised speech representation. It is expected that leveraging their benefits synergistically in a unified solution would further improve the overall system performance. This paper presents a novel design of a hierarchical information path, named LUPET, which sequentially encodes, from the shallow layers to deep layers, multiple aspects of linguistic and acoustic information at diverse granularity scales. The path starts from LID prediction, followed by acoustic unit discovery, phoneme sharing, and finally token recognition routed by a mixture-of-expert. ASR experiments are carried out on 10 languages in the Common Voice corpus. The results demonstrate the superior performance of LUPET as compared to the baseline systems. Most importantly, LUPET effectively mitigates the issue of performance compromise of high-resource languages with low-resource ones in the multilingual setting.
keywords:
Multilingual ASR, language identity, self-supervised speech representation learning, mixture-of-expert
@@ 1 Introduction
Conventionally an automatic speech recognition (ASR) system is developed to transcribe speech into text for a specific language. Toward multilingual ASR, recent research is focused on building a unified model that covers multiple languages [1, 2, 3, 4]. One practical advantage is the reduction of training and deployment costs, as compared to building a separate monolingual model for each language. It also facilitates sharing of linguistic knowledge among the languages and may help elevate the recognition performance on those with limited data resources [5]. It was shown that training a fully shared end-to-end (E2E) model on a multilingual speech corpus is a simple yet effective solution (vanilla[4]. The multilingual corpus is made by mixing the corpora of different languages, and a shared vocabulary is used. Due to the heterogeneous nature of different languages [6, 4], the vanilla scheme exhibits the issue that the recognition performance on the high-resource languages inevitably compromises in the multilingual training, in order to attain reasonable performance on the low-resource languages.
To mitigate the issue of performance compromise, there have been attempts that incorporate language identity (LID) information [7, 8, 9, 10], phoneme information [11, 12, 13], and language-specific architecture, e.g., mixture-of-expert (MoE) [14, 15, 16, 17, 18]. These approaches typically require a supervised training process, which requires labeled training data. On the other hand, self-supervised learning (SSL) is believed to be an effective way of cross-lingual data sharing. The representative works include wav2vec2.0 [19], HuBERT [20], XLSR [21], and many others. The two-stage training scheme, i.e., pre-training and fine-tuning, has been widely applied in SSL. In [22], joint unsupervised and supervised training (JUST) was shown to outperform two-stage training on multilingual ASR. JUST uses a contrastive loss and a masked language model (MLM) loss to learn discrete units for better contextualized representations.
In the present study, a hierarchical information path is developed to combine multiple useful factors synergistically to boost the overall performance of a multilingual ASR system. The path comprises a sequence of prediction modules that incorporate linguistic and acoustic information at diverse granularity levels into the recognition process. These modules are namely, LID, Acoustic Unit discovery, Phoneme sharing, and mixture of Experts for Token recognition. We use the acronym LUPET to denote the proposed design, in which each alphabet represents one of the information components in the path. The path LUPET can be easily integrated into a vanilla ASR architecture by unfolding with the encoder layers. Within this path, information from a shallow layer can benefit those that occur in deeper ones, and hence it is considered a hierarchical flow. Here the shallow layer refers to an encoder layer close to the input. Importantly, the required information labels are either straightforward to obtain or can be derived via an SSL process.
The effectiveness of LUPET is evaluated by experiments on 10 languages in the Common Voice [23]. The results show that, compared to the vanilla system, LUPET can achieve 19.7% and 12.3% relative reduction of average word error rate with CTC and attention decoding, respectively. LUPET also outperforms previous baseline systems. In particular, it demonstrates superior performance on the high-resource languages as its performance compromise to low-resource languages is alleviated.
@@
2 LUPET
@@ 2.1 Vanilla E2E Multilingual ASR
The E2E multilingual ASR architecture we adopt is the hybrid CTC-Attention conformer [24, 25]. It consists of three components, namely encoder, decoder, and CTC [26] layer. The encoder takes an acoustic feature sequence $\mathbf{X}=\{\mathbf{x}_{t}\}_{t=1}^{T}$ as input and converts it to hidden representation $\mathbf{H}=\{\mathbf{h}_{t}\}_{t=1}^{T^{{}^{\prime}}}$, where $T$ and $T^{{}^{\prime}}$ denote the number of original frames and the sub-sampled frames. The $\mathbf{H}$ is then forwarded to two classification branches for predicting the token sequence $\mathbf{Y}=\{y_{u}\in\mathcal{V}\}_{u=1}^{U}$, where $\mathcal{V}$ is a shared multilingual vocabulary built by BPE [27] and $U$ denotes the number of tokens. One classification branch, i.e., decoder, conditions $\mathbf{H}$ to autoregressively compute token-level posterior $p(y_{u}|\mathbf{H},y_{1:u-1})$ via a cross attention mechanism. The attention loss is given as
$\vspace{-2pt}\mathcal{L}_{Attn}=-\sum_{u=1}^{U}logp(y_{u}|\mathbf{H},y_{1:u-1}).\vspace{-2mm}$ (1)
@@
where $\mathbf{z}_{t}$ is the logits over $\mathcal{V}\cup\emptyset$ and $B^{-1}$ is the inverse function that gives all valid alignment paths between input sequence $\mathbf{H}$ and output sequence $\mathbf{Y}$. The blank token $\emptyset$ is specially designed by CTC for aligning $\mathbf{H}$ and $\mathbf{Y}$. $\mathbf{Z}=\{\mathbf{z}_{t}\}_{t=1}^{T^{{}^{\prime}}}$ represents one possible alignment path. The training objective of hybrid CTC-Attention is a linear combination of Eq. 1 and Eq. 2:
@@ 2.2 Incorporating LUPET
Recent studies have pointed out (1) Both LID and phoneme information are beneficial for multilingual training [10, 13]; (2) Design language-specific modules to process language-specific information is useful to reduce language interference [16, 18]; (3) The success of self-supervised cross-lingual representation learning applied in ASR [21, 28]. Although these factors’ effectiveness has been verified separately, how to synergistically combine them to contribute a better solution from a unified perspective remains an open question. The proposed LUPET provides a novel view from the multilingual hierarchical information path. From LID to acoustic unit followed by phoneme then go through MoE routing to the final token, the information that occurred in the early position of the path is assumed to contribute to the prediction of later information of the path.
@@
where $LIN$ denotes a linear layer to keep the hidden dimension and $\mathbf{H}^{s^{\prime}}$ is the input representations of $Enc^{lm}$.
The lower-middle layers of encoder $Enc^{lm}$ are utilized to perform acoustic unit discovery. Similar to BEST-RQ [28], a random-projection quantizer including a projection matrix $Proj^{c}$ and codebook $\mathbb{C}$ is applied and none of the parameters are trainable. Vector quantization (VQ) is carried out on acoustic features $\mathbf{X}$ to produce discrete labels
@@
where $Sub$ represents the subsample operation and $Proj^{c}$ performs projection from the speech feature dimension to the code vector dimension. The output of $\mathbb{C}$ are the indices of the nearest code vectors of the codebook to the input vectors.
@@
To incorporate the hierarchical information, from LID, acoustic unit, phoneme, and token, the objective function of LUPET is given as a linear combination of Eq. (3, 4, 7, 8):
@@ 3.1 Dataset
The 10 languages, namely English (en), French (fr), Spanish (es), Chinese (zh), Italian (it), Russian (ru), Portuguese (pt), Turkish (tr), Dutch (nl) and Tatar (tt) from the public available Common Voice 13.0 [23] are selected for our multilingual ASR experiments. The language coverage includes high-resource languages, e.g., English with around 2,280 hours of training data, and low-resource languages, e.g., Tatar with only about 20 hours. The detailed training and testing statistics are listed in Tab. 1. Note that zh includes Mandarin, Taiwanese, and Cantonese. A standard text normalization (the same as in Whisper [2]) is applied to all transcriptions of the dataset.
@@ 3.2.1 Vanilla
The vanilla model adopts a hybrid CTC-Attention architecture. The encoder has 12 conformer layers with 8 attention heads and 512 hidden dimensions, while the decoder has 6 transformer layers [29]. The CTC weight $\lambda$ in Eq. 3 is set to $0.3$. The input acoustic feature to the network is the typical 80-dimensional log-Mel filterbank. The output vocabulary used is derived from Whisper’s tokenizer. This tokenizer was obtained by BPE using UTF-8 bytes of the entire training dataset of Whisper.
@@ 3.2.2 LUPET
Compared to vanilla, the encoder architecture has several modifications by incorporating LUPET. The output positions of { $Enc^{s}$, $Enc^{lm}$, $Enc^{um}$, $Enc^{d}$} are at the {3-th,6-th,9-th,12-th} layer of the original encoder, respectively. When MLM loss takes effect, acoustic feature $\mathbf{X}$ is randomly masked consecutive 20 frames with probability $p=0.01$. Codebook $\mathbb{C}$ of the random-projection quantizer has a size of 8192 and a dimension of 16. IPA sequence per-utterance is obtained using an open-sourced toolkit phonemizer [30]. In each layer of $Enc^{d}$, MoE including 8 FFN experts is used to replace the end-FFN of the original conformer layer. In Eq. 10, the weight coefficients $w_{1}$, $w_{2}$, and $w_{3}$ are set to 0.3, 0.07, and 0.3, respectively.
@@ 3.2.3 Baselines
Several baselines are used for comparison: (1) Mono, monolingual ASR with vanilla architecture and 256 hidden dimensions is trained per language. (2) Oracle_LID, append the pre-known LID embedding to input acoustic feature for multilingual ASR training. (3) MoE [16], keep the $Enc^{d}$ of LUPET and remove other auxiliary losses, hidden representation $\mathbf{H}^{um}$ is used as the input to the routing network. (4) LID_SC [10], LID prediction by CTC and LID information self-conditioning (SC) are performed over the vanilla model. (5) Whisper, whisper-large-v2 with oracle LID is used for decoding.
@@ 3.3 Training Scheme and Evaluation Metric
We implement the vanilla and our proposed LUPET methods on the Wenet toolkit [31]. The model is trained with Adam [32] optimizer with a learning rate (LR) of $31e-$ 3. LR schedule has a warmup step of $15000$. Batch size is set to $12$ with $accum\_grad=16$. 8 V100 GPUs are used for DDP training. Each multilingual model is trained for 50 epochs and each monolingual model is trained for 100 epochs. If not specified otherwise, MLM takes effect from epoch 5 to 30. The final model for decoding is obtained by averaging the $10$ best models with the lowest validation losses. Character error rate (CER) for Chinese and word error rate (WER) for other languages are used to measure the system’s performance.
@@ 4 Results and Analysis
Table 2: WER (%) results of different systems on 10 languages of Common Voice by CTC greedy decoding. avg 5high denotes the averaged WER results of top-5 high-resourced languages and avg 5low is similar for the low-resourced case. In the LUPET block, the backslash /  represents the ablation study by removing the following component, where U = acoustic unit discovery, P = IPA sharing, and L = LID prediction. $w_{2}=1$ means the weight coefficient of $\mathcal{L}_{mlm}$ is set as 1. Uto50ep means U takes effect until 50 epochs.
@@ 4.1 Performance Comparison to Monolingual System
A desirable multilingual ASR is expected to present better performance than its monolingual counterparts (Mono). As shown in Fig. 2, LUPET and other baselines are used to compare performance to monolingual systems on the test sets of 10 languages. The x-axis follows an order from high-resource to low-resource languages. The y-axis denotes the relative WER to Mono, where the more negative value represents the lower WER. It is clear to see all curves are basically decreasing except the peak at ru. The decreasing trend is straightforward as the more low-resourced languages can achieve more significant performance gains. Multilingual training brings degradation to the Russian (ru) language in most cases. We speculate this may be due to the compromised phenomenon from Russian (ru) to Tarta (tt). The language tt achieves above 60% relative WER reduction via multilingual training at the cost of side effects on language ru, belonging to a different language family with tt.
@@
4.2 LUPET’s Effectiveness Verification
Tab. 2 presents the WER results of different systems on 10 languages. Averaged WER are calculated for overall comparison. The top-5 languages, with more training data, are roughly referred to as high-resources (5high), while the remaining 5 languages are low-resources (5low). Having similar observation from Fig. 2, LUPET gives significantly better performance on all languages compared to other baselines. The benefits of LID prediction and MoE routing structure have been well verified by system LID_SC and MoE.
@@ 4.2 LUPET’s Effectiveness Verification
Disabling both U and P (IPA sharing prediction) exhibits worse results. Tab. 2 provides two comparison views to understand the independent effect of the component P. (1) LUPET / LU can be seen as “MoE + P”. Comparing it with MoE, IPA sharing is found to be beneficial for all languages except tt. (2) When only removing P from LUPET, WERs on all languages basically degrade. Low-resource languages clearly give the worse results, where tr increase the absolute WER above 4%. One possible reason is that the independent U would lead to information loss due to the masking mechanism in MLM, especially for low-resource languages. In LUPET, with the help of intermediate phoneme prediction (P), the loss of information is largely mitigated, thus not presenting much worse results on low-resource languages.